Fermi National Laboratory
 

Accelerator Update for March 26, 2007 - March 28, 2007

 

Monday March 26

The day shift began with Operations monitoring Tevatron (TeV) store 5300 (156.89E30) the Recycler stash (58.8E10), with the Antiproton Source (Pbar) stacking (40.3E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Pbar experts began a study at 9:23 AM. There was no stacking during the study.

The Pbar study ended and stacking resumed at 10:11 AM.

At 10:27 AM, MI beam aborted due to bad beam positions for the first of four times in four hours. It reset okay.

The Neutron Therapy Facility (NTF) reported at 11:25 AM that they had treated their last patient for the day.

At 11:47 AM, the MI beam permit tripped due to the MI-8 collimators. Everything reset.

At 12:48 PM, a Linac RF station (LRF4) tripped off. Experts investigated and replaced a keyer driver. Linac resumed sending beam at 1:41 PM.

At 2:16 PM, a Tevatron Electron Lens (TEL1) had its field drop from 37 kG to 17.8kG. Operations contacted experts.

At 3:58 PM, LRF4 tripped off for the first of two times in 13 minutes. Operators reset it locally each time.

Operations reported at 4:02 PM that the Recycler had stashed antiprotons twice during the day shift.

At 4:28 PM, LRF4 tripped off again. Operators reset it locally.

At 5:23 PM, Operations began preparing the Booster for an access. RF experts will repair BRF10. Linac experts will work on LRF4 during the access.

Linac experts completed their work on LRF4 at 7:10 PM.

RF experts completed their work on BRF10 and were out of the tunnel by 7:36 PM.

Linac resumed sending beam at 9:33 PM.

At 9:33 PM, LRF4 tripped off for the first of two times in one minute. It reset.

At 11:34 PM, the NuMI beam permit tripped off due to a trim power supply (V118). Operators investigated and found the MI-65 power supply room hot. Operations informed the Duty Mechanic about the room temperature. The power supply would not reset. A power supply expert was called in.

Tuesday March 27

The midnight shift began with Operations monitoring store 5300 (51.73E30) and the stash (173.5E10), with Pbar stacking (8.0E10), and with NuMI off due to a E:V118 problem.

At 12:37 AM, the MI beam permit tripped for the first of 13 times in about two hours. It reset each time.

A power supply expert reported that he had completed his repairs on V118. Operations resumed sending beam to NuMI at 2:37 AM.

The Duty Mechanic reported at 3:47 AM that he had replaced a broken belt for the MI-65 power supply room’s air conditioner.

At 4:31 AM, Operations reported a problem with the Pbar target rotation. Operators tried to reverse the target’s rotation, but it would not move. Operations contacted an expert.

At 5:12 AM, NuMI tripped off due to a MI-52 kicker coolant temperature alarm error. It reset.

At 5:27 AM, NuMI tripped off on coolant temperature for the first of six times in about 45 minutes. Operations reduced NuMI’s beam intensity and this brought the temperature down.

Operations established beam to MTest at 6:37 AM.

Operations prepared Pbar for an access at 7:25 AM to repair the target motor.

At 7:28 AM, a TeV RF station (TRF4) tripped off. It caused a loss of luminosity.

At 7:30 AM, the 750keV area buncher tripped off. Operators reset it locally.

At 8:04 AM, LRF3 tripped off for the first of three times in two minutes.

Operations reported 8:05 AM that the Recycler had stashed twice during the midnight shift.

Operations turned off the Booster beam at 10:01 AM for MI-65 transport line safety system tests.

A NuMI expert accessed MI-52 at 10:02 AM and cleaned the kicker coolant strainer.

The MI-65 transport safety system testing finished at 10:51 AM.

The NuMI kicker work ended and operations established beam to NuMI at 11:05 AM.

At 11:27 AM, LRF4 tripped off. It reset.

Operations terminated store 5300 at 11:34 AM. Experts began work on TEL1.

Operations put the TeV into shot setup at 11:50 AM.

At 12:05 PM, LRF4 tripped off. It reset.

The work on the Pbar target motor was completed by 12:10 PM.

Experts completed their TEL1 work at 12:15 PM.

AT 1:18 PM, Operations reported a Pbar power supply (H717) LCW leak. Experts began repairs.

Power supply experts reported at 2:38 PM that they had completed the repair of the Pbar H717 LCW leak.

At 2:48 PM, BRF18 tripped off. It reset.

At 2:48 PM, Linac experts turned off LRF4 to make adjustments. The station’s driver tripped off due to a shorted screen. They cleared the short. Linac resumed sending beam 30 minutes later.

Recycler experts began a study period at 3:05 PM.

Operations established store 5301 at 3:06 PM with an initial luminosity of 204.41E30.

At 3:41 PM, experts turned off the NuMI horn for about 10 minutes to repair a LCW leak in its backup power supply.

At 3:55 PM, Operations reported to TeV experts that the TEL1 power supply had a LCW leak.

Operations reported at 4:44 PM that the TEL1 LCW power supply leak was repaired.

At 4:44 PM, LRF3 tripped off. It reset.

At 4:45 PM, BRF18 tripped off. It reset.

At 5:20 PM, the Recycler suffered from a vacuum burst. Many vacuum pumps tripped off. Everything reset okay.

At 5:53 PM, LRF2 tripped off. It reset.

The MTest T950 and T964 experiments reported that they were through with the beam at 6:08 PM. Their run had ended. A new experiment (T965) will begin running on Thursday afternoon or Friday morning. T965 will test a newly designed photosensor, a Pixelated Silicon Photosensor (PSiPs), for use as a readout of scintillator and WLS-fiber detectors.

Operators prepared the Meson M01-5 enclosures for access at 6:11 PM.

Recycler studies ended at 7 PM.

At 9:09 PM, LRF4 tripped off. It reset.

Operations reported 11:54 PM that the Recycler had stashed twice during the evening shift.

Wednesday March 28

The midnight shift began with Operations monitoring store 5301 (75.28E30) and the stash (90.1E10), with Pbar stacking (51.2E10), and with NuMI taking beam.

At 2:32 AM, LRF5 tripped off. It reset.

At 2:43 AM, the Pelletron tripped off. Operators restored a file and reset the Pelletron.

At 3:13 AM, Operations reported the NuMI beam permit had tripped off several times due to beam positions. NuMI tripped off a number of times more `over the next four hours due to the same problem.

At 5:56 AM, a Booster vacuum pump tripped off and brought down BRF13 &14. Everything reset okay.

At 7:34 AM, Operations reported problems with BRF17 & 18. Operators lowered the beam intensity, turned off and bypassed BRF17 & 18, and turned on BRF19.

Operations reported 8:19 AM that the Recycler had stashed four times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar experts to conduct studies this morning and has requested MI and Booster experts to work on beam quality issues.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/28/2007    email Fermilab